If a camera with a 60 mm lens is to resolve two objects that are 3.0 mm from each other and both objects are 4.0 m from the camera lens, what is the minimum diameter of the camera lens? Assume the wavelength of the light is 550 nm

Express your answer using two significant figures

What is the revolving power?

Express your answer using two significant figures.
Answer #1

focal length f = 60mm = 0.006 m

seperation between two objects d = 3 mm = 0.003 m

distance from camera L = 4m

lambda = 550 nm = 550 x 10^-9 m

Using Rayleigh Criterian

d/L = 1.22* lambda / D

D = 1.22 x 550 x 10^-9 * 4 / 0.003 = 0.89 mm


Resolving power

s = 1.22 * lambda * f / D = 1.22 * 550 x 10^-9 * 0.006 / 0.00089 = 4523.6 x 10^-9

s = 4.524 x 10^-6 m